### v2

# Gyromorphs: a new class of functional disordered materials

Abstract
We introduce a new class of functional correlated disordered materials, termed Gyromorphs, which uniquely combine liquid-like translational disorder with quasi-long-range rotational order, induced by a ring of $G$ delta peaks in their structure factor. We generate gyromorphs in $2d$ and $3d$ by spectral optimization methods, verifying that they display strong discrete rotational order but no long-range translational order, while maintaining rotational isotropy at short range for sufficiently large $G$. We show that they are distinct from quasicrystals, which we argue are akin to Fourier “duals” of gyromorphs. Using a coupled dipoles approximation, we show that these structures outperform quasicrystals, stealthy hyperuniformity, and Vogel spirals in the formation of low-index-contrast isotropic bandgaps in $2d$, for both scalar and vector waves, and open complete isotropic bandgaps in $3d$. Finally, we introduce “polygyromorphs” with several rotational symmetries at different length scales (i.e., multiple rings of delta peaks), enabling the formation of multiple bandgaps in a single structure, thereby paving the way for fine control over optical properties.
preprint: APS/123-QED
A crucial aspect of the design and discovery of new materials is understanding the relationship between structure and properties. Crystals have proven to be a highly versatile platform for engineering functions, as the periodicity of their atomic arrangement greatly facilitates the prediction and optimization of their properties. However, not all properties can be realized with periodic structures. Aperiodic media can achieve transport properties unattainable in periodic systems, such as the formation of isotropic photonic bandgaps, which are highly desirable in optoelectronic applications such as freeform waveguides [1, 2] and tunable-reflectance coatings [3]. While isotropic bandgaps have been demonstrated in some deterministic aperiodic systems such as Vogel spirals [4, 5, 6, 7, 8], and near-isotropic bandgaps have been observed in quasicrystals [9, 10, 11], many recent works have sought this property in correlated disordered structures [12, 13], i.e. random point patterns that lack conventional long-range order but exhibit spatial correlations.
Unlike in periodic systems, the structural origins of photonic bandgaps – or more precisely, “pseudogaps”, characterized by a depletion in the density of states relative to vacuum, leading to reduced transmission [14, 15] – remain poorly understood in aperiodic systems. Disordered stealthy hyperuniform (SHU) structures [16], which suppress density fluctuations over long distances, have been shown to exhibit isotropic bandgaps in both $2d$ and $3d$ [17, 18, 19, 20, 14, 21, 15, 22], but only when strong short-range correlations are also present [17, 23, 14, 24, 15, 25]. Further complicating the picture, photonic bandgaps have also been observed in non-SHU disordered structures, such as jammed packings [26, 27, 28], and even in systems with only short-range correlations, like equilibrium hard spheres [29], which are not hyperuniform. The role of short range correlations in disordered systems is especially pronounced for vector waves, for which bandgaps have so far only been observed in networked materials, such as honeycomb or tetrahedrally bonded structures  [10, 17, 27, 30, 31]. In these random networks, the emergence of bandgaps has been linked to the similarity of local geometry and topology across the network [23, 27, 30].
Figure 1: Introducing: Gyromorphs. $(a)$ Section of the point pattern of a $60$-fold gyromorph. We show the shortest distance displaying 60-fold order. $(b)$ Corresponding structure factor. $(c)$ Corresponding pair correlation function $g(\bm{r})$ near the origin. $(d)$ Radial structure factor $S(k)$ (top) and radial distribution function $g(r)$ (bottom) for the 60-fold gyromorph.
This raises the question: is there a unifying feature across these systems that promotes the formation of isotropic photonic bandgaps? In the single-scattering regime, the emergence of a bandgap can be attributed to the presence of strong scattering at some characteristic frequency, and of weaker scattering at neighboring frequencies. This argument has been invoked to justify the appearance of bandgaps in quasicrystals [32, 33, 34, 11], and explains why bandgaps also arise in systems with short-range correlations, such as equilibrium hard sphere liquids [29, 35]. In fact, we observe that all previously cited aperiodic systems that exhibit isotropic bandgaps have an isotropic ring of high values in their structure factor, $S=|\hat{\rho}(\bm{k})|^{2}/N$, where $\rho(\bm{r})$ is the density field defining the system. This condition is also well approximated by quasicrystals, whose most intense structure factor peaks form regular polygons [36, [SeeSupplementalMaterialat][foracomparisonbetweenstructurefactorsofknownbandgapformersandgyromorphs, detailednumericalmethodsincludingacharacterizationofthenatureandvarietyofstructuresreachedbyouralgorithm, measurementsoftheeffectofrandomkicksongyromorphs, furtherdiscussionsaboutthedifferencesbetweengyromorphsandquasicrystals, abriefdescriptionofthesumofcosinefunctionthatapproximates\$g(r)\$, areminderofthederivationofrelevantquantitiesandequationswithinthecoupleddipolesmethod, additionalscalingsforthebandgapqualityagainstphysicalparameters, theequivalentofthemaintextresultsfor\$2d\$and\$3d\$scalarwaves, andashortcomparisonbetweencoupleddipolesandfinite-differencemethods.]supp].
Bravely adhering to this single-scattering rationale, an ideal bandgap material should exhibit a ring of high values in $S(k)$, contrasting with low values around it. Accordingly, we propose a correlated disordered structure of non-overlapping points whose structure factor displays one ring of $G\in 2\mathbb{N}^{\star}$ delta peaks with intensities $\mathcal{O}(N)$ on a circle with radius $K$, but as little order as possible elsewhere. In this work, we show that such hypothetical structures, which we call gyromorphs, are realizable, and we present a reliable algorithm for their generation in $2d$ and $3d$.
In gyromorphs, liquid-like isotropic neighborhoods coexist with a ring of extensive $S(k)$ peaks, resulting in quasi-long-range rotational order. These structures are thus fundamentally distinct from any previously known materials, as they reconcile seemingly contradictory features. We successfully generate finite gyromorphs with up to $G\sim 10^{3}$ peaks, whose heights surpass the peak intensities of finite quasicrystals obtained by usual deterministic methods [38, 39] (see Appendix). Remarkably, we uncover a duality between the structure factor and pair correlation functions of gyromorphs and quasicrystals, suggesting that, at high rotational symmetries, gyromorphs may outperform quasicrystals as bandgap materials. Using a coupled dipoles approximation [40, 41, 42], we thus demonstrate that gyromorphs outmatch previous candidate systems in $2d$ as low-index isotropic bandgap materials, for both vector and scalar waves [37]. Finally, we demonstrate that gyromorphs can be extended to $3d$ structures exhibiting complete isotropic bandgaps, as well as to “polygyromorphs” with multiple rings of peaks, resulting in multiple bandgaps in a single system.
Generating gyromorphs – To generate gyromorphs, we use the Fast Reciprocal-Space Correlator (FReSCo) [35]. We use the NUwNU (non-uniform real space with non-uniform k-space constraints) variant, which imposes constraints at continuous reciprocal space positions, assuming free boundary conditions for the point pattern. Starting from an initially uncorrelated random point pattern with $N$ points, we optimize it to display $G$ $\mathcal{O}(N)$-high peaks regularly spaced on a ring while forbidding overlaps, thereby creating a $G$-fold gyromorph (see SM [37]).
Figure 2: Order in gyromorphs. $(a)$ $S(k)$ of a $60$-fold gyromorph (blue) with $N\sim 10^{5}$ ($KL/2\pi=300$), compared to a Percus-Yevick $S(k)$ for $\phi=0.57$ (dashed red line) in log scales. $(b)$ Corresponding radially averaged profile of peaks across $K$ values, rescaled by peak height (colored lines) a solid black line indicates the radial profile of $\text{sinc}^{2}(k_{x}L/2)\text{sinc}^{2}(k_{y}L/2)$. Inset: Rescaled peak height $GS(\bm{k}_{p})$ against $N$, in log scales, across $G$ (colored symbols). A dashed black line shows $GS(\bm{k}_{p})=3.5N$.
Structure of gyromorphs – At high $G$, although the point pattern appears isotropic at short length scales, gyromorphs still exhibit strong $G$-fold order at large length scales, as shown in Fig. 1. In Fig. 1$(a)$, we show a portion of an $N\sim 10^{4}$, $G=60$ gyromorph, highlighting in red the “gear” with radius $R_{G}\approx G/K$ [37] onto which $60$-fold order is first achieved around an example point. Note that this is different from bond-orientational order [43], which describes nearest-neighbor ordering only; at high $G$, gyromorphs have no such order, unlike quasicrystals [37]. In Fig. 1$(b)$, we show the corresponding structure factor, highlighting a strong $60$-fold ring of peaks, that are intense enough to display a few echoes, and low-intensity regions around the peaks. In Fig. 1$(c)$, we show the central part of the corresponding pair correlation function $g(\bm{r})$. This function illustrates that each particle is surrounded by the gear of Fig. 1$(a)$, with the same orientation, at the smallest ring of neighbors capable of accommodating $G$-fold order [37]. Additionally, it reveals that, at shorter distances, the local neighborhoods are perfectly isotropic. The short-range “disorder” is highlighted in Fig. 1$(d)$ by the radial plots of $S(k)$ (top) and $g(r)$ (bottom). While $S$ is peaked at $K$ and shows a few nearby echoes, it eventually decays to $1$ at larger $k$, indicating the absence of short-range order. Similarly, $g$ is notably less peaked than that of (quasi)crystalline structures, exhibiting no visible feature at $R_{G}$, and decaying with increasing distance. Overall, Fig. 1 highlights that $S-1$ is essentially just a $G$-fold ring of Dirac deltas, so that the $g(r)$ is well approximated by a sum of cosines akin to “kaleidoscopic” optical fields obtained by holographic techniques [44, 45]. In contrast, quasicrystals display rings of peaked $g$ corresponding to a discrete set of allowed neighbor positions [38, 39], so that $S$ is well approximated by a sum of cosines at high $G$. This “Fourier duality” between gyromorphs and quasicrystals is illustrated in Appendix.
To further investigate the nature of order in gyromorphs, we replot a structure factor in log-log scales in Fig. 2$(a)$. This panel shows that gyromorphs are not hyperuniform, as the low-$k$ limit of $S$ saturates at a level similar to that of a Percus-Yevick model for equilibrium hard disk liquids [46]. We further validate this observation by measuring local number fluctuations in real space over windows of increasing size [16, 35] (see Appendix), finding that the long-range density fluctuation scaling of gyromorphs is indistinguishable from that of hard disks. Furthermore, at large $k$, the decay of $S(k)$ is also similar to that of a liquid. These results suggest that, with respect to translational order, gyromorphs are more akin to liquids than to quasicrystals.
In Fig. 2$(b)$, we analyze the peaks. In the main panel, we show the profile of the peak at $(KL/2\pi,0)$, normalized by its height, as a function of the distance $kL/2\pi$ to the maximum, across system sizes for $G=60$. As $N$ grows, the profile converges to $\textrm{sinc}^{2}(k_{x}L/2)\textrm{sinc}^{2}(k_{y}L/2)$, implying that the linear width of the peaks in $k$-space decays like $1/L$. The inset shows that, for different values of $G$, the peak heights grow like $\mathcal{O}(GN)$, indicating that the peaks are extensive for any fixed order symmetry. Altogether, the peaks have extensive height while their area decays like $1/L^{2}\sim 1/N$: they approach Dirac deltas as $N$ increases, like in quasicrystals, so that gyromorphs display quasi-long-range rotational order (see Appendix).
Coupled Dipoles Method – We now turn our attention to the optical properties of gyromorphs. From an Ewald-sphere construction [47, 48] one expects a ring of many high peaks at radius $K$ to generate strong backscattering at $K/2$ [33]. That is why quasicrystals were deemed good candidates for isotropic bandgaps. However, they display increasingly shallow bandgaps as their order of rotational symmetry grows [10], which can be attributed to weakening features in $S$ in de Bruijn’s construction [38, 39] as $G$ increases. By contrast, gyromorphs are only constrained to display a finite set of high fine peaks, and they may retain a deep isotropic bandgap even at high $G$.
Figure 3: Optical properties of gyromorphs We use an example 60-fold gyromorph and focus on TE polarization. $(a)$ Intensity transmission for frequencies $40\leq k_{0}L/2\pi\leq 60$ (radial direction) and $360$ incident angles (orthoradial direction) of a source Gaussian beam. $(b)$ Relative LDOS change with respect to vacuum, averaged over $1000$ random points, highlighting a dip close to $k_{0}=K/2$ (here $KL/2\pi=100$). $(c)$ Spatial map of the intensity at angle 00 and $k_{0}L/2\pi=47.6$, which corresponds to the minimum of both transmission and LDOS. $(d)$ Corresponding LDOS map.
To verify this, we assume that each point in the gyromorph represents a non-magnetic scatterer with a dielectric contrast $\delta\varepsilon(\omega)\in\mathbb{C}$ from the surrounding medium at pulsation $\omega$, corresponding to a refractive index $n=\sqrt{1+\delta\varepsilon}$, and a radius $a$. Limiting our analysis to wave-vectors $k_{0}=\omega/c$ (with $c$ the celerity of waves in the host medium) such that $k_{0}a\ll 1$ (Rayleigh scattering [42]), we employ a coupled dipoles approximation to describe the scattering of waves by the structure. This standard method [40, 41, 42] turns the problem of solving Maxwell’s equations in the medium (a very strenuous task that involves a full discretization of space as well as very specific choices of boundary conditions [49, 50]) into a linear system, see Appendix. In short, setting only $a$ and $n$ for the scatterers, we measure transmitted intensities for any choice of monochromatic source field, as well as an integrated local density of optical states (LDOS). Throughout the main text, we present results for vector waves (Transverse Electric, or TE, mode in $2d$, as opposed to the scalar-wave Transverse Magnetic, or TM, mode). We report the scalar-wave results in SM [37], as scalar-wave bandgaps are less challenging to obtain in point patterns [10, 17, 23, 15]. All optical properties are computed within the Materials Analysis via Green’s Tensors (MAGreeTe) library, which we make publicly available [51].
Existence of a bandgap – First, in Fig. 3, we focus on a single $2d$ gyromorph with $G=60$ and $N\sim 10^{4}$ to determine whether a bandgap is present. We first trim the system to a circular shape of radius $R=L/2$ to avoid any anisotropy due to the overall shape of the medium, like in many experimental tests [9, 17, 18]. We then set the radius $a$ of scatterers such that the filling fraction $\phi\equiv Na^{2}/R^{2}$ equals $5\%$ and the dielectric contrast $\delta\varepsilon=8$ ($n=3$ for scatterers in vacuum), a typical value for metal oxides used in experiments [22]. Following Refs. [9, 17, 18, 22], we first perform transmission measurements. We set the source to be a standard Gaussian beam with zero curvature focused at the center of the medium [52, 53, 42]. We solve the coupled dipoles system for $360$ regularly spaced orientations $\theta$ of the beam and for several frequencies $k_{0}$. Then, following Refs. [52], we measure the transmitted intensity behind the system at a distance $D$ from its nearest edge  [^1] The reason for not measuring the transmission in the far field is that the monochromatic Gaussian beam is not in fact a physical laser beam, in the sense that its intensity does not decay according to the Rayleigh-Sommerfeld boundary condition in the far field [42].. For each incoming beam orientation $\theta$, we define the transmission as the average intensity over 180 angles in $[\theta-\pi/2;\theta+\pi/2]$ on the half-circle with radius $D$, normalized by the value of the incident field at these points in the absence of the medium. The results, shown in Fig. 3$(a)$ in the plane of incident wave-vectors $(k_{0},\theta)$, display a clear trough of low intensities at $k_{gap}\lesssim K/2$, as expected from single-scattering arguments [33, 35]. An intensity map for $\theta=0$ and $k_{0}=k_{gap}$ is shown in Fig. 3$(c)$, confirming that the field is strongly backscattered. To confirm that this lower transmission is the sign of a bandgap, we measure the DOS of the system. In Fig. 3$(b)$, we show $\delta\varrho$, the relative change of DOS compared to vacuum averaged over $1000$ random points drawn randomly across the material but at least $2a$ away from scatterers, following Refs. [55, 14]. We show strong mode depletion at $k_{gap}L/2\pi=47.6$. Finally, a map of the local $\delta\varrho$ across the system in Fig. 3$(d)$ shows consistent mode depletion across the bulk. We show similar TM results in SM [37].
Figure 4: Comparison with other systems Top row: TE intensity transmissions as a function of $k_{0}$ (radial) and beam orientation (orthoradial) for (from left to right) $10$- and $60$-fold de Bruijn quasicrystals, an SHU structure, and a Vogel spiral. Bottom row: Corresponding $\delta\varrho$ (red lines), with the $60$-fold gyromorph shown as a light gray line. All results are obtained for $n=3$, $\phi=0.05$.
Comparison to other systems – We compare the results of Fig. 3 to classic bandgap candidates, namely $10$- and $60$-fold deterministic quasicrystals obtained by the de Bruijn’s construction [38, 39, 56], a deterministic Vogel spiral [4], and a random SHU structure generated by FReSCo [35] with a degree of stealthiness such that it displays a TM bandgap [17, 18, 29, 14, 15, 21]. Structures are adjusted to contain $N\sim 10^{4}$ points and a ring of higher values at the same $K$ as the gyromorphs, and we set all parameters to the same values as in Fig. 3 (see SM [37] for $S(k)$). We report transmission plots in the top row of Fig. 4, and DOS in the bottom row. We show that Vogel and SHU systems display much shallower transmission troughs than the gyromorphs, which we attribute to weaker features in $S(k)$, and very little DOS change. Furthermore, we show that quasicrystals may either display rather deep but anisotropic gaps at low symmetries, or isotropic but shallow gaps at high symmetry due to the broadening of their $S(k)$ features [10, 37]. In fact, the bandgaps of finite-size gyromorphs are as deep as in low-order quasicrystals but more isotropic than them at comparable $\delta\varepsilon$. We check [37] that the DOS dip in gyromorphs gets deeper with index contrast and system size (following a power law, like in defective crystals [57, 58]), but does not strongly depend on $G$.
Figure 5: Three-dimensional gyromorphs. $(a)$ 3d-rendered orthographic view of the constrained peaks on the sphere. From an icosahedron (red), we subdivide edges (blue) and faces (green) so as to cover the whole sphere uniformly. $(b)$ Cross-sections (in the $x=0$, $y=0$, and $z=0$ planes) of the $g(\bm{r})$ of a gyromorph with $G=1212$, $N=83762$. $(c)$ Vector-wave transmission averaged over the forward half-sphere, for a beam source rotating in the $z=0$ plane. $(d)$ Relative DOS change for the gyromorph (red line), compared to that of a Poisson point pattern (gray line). In $(c),(d)$, we use a ball of $N\approx 15000$ points from the structure and set $\phi=0.1$, $n=6.5$.
Extension to $3d$ We now extend the concept of gyromorphs to $3d$ structures. We optimize structures to display $G$ regularly spaced peaks on a sphere with radius $K$ in their structure factor. As there is no standardized way to populate points evenly spaced on a sphere while preserving centrosymmetry, we choose the location of peaks by building upon icosahedral geometry. Starting from an icosahedron of peaks, we subdivide each icosahedral face into multiple triangular faces and project those additional vertices onto the sphere. An example of such a set of peaks is shown in Fig. 5$(a)$. We show in Fig. 5$(b)$ cross-sections of the $g(\bm{r})$ for a gyromorph thus obtained with $G\approx 10^{3}$. Finally, in Fig. 5$(c)-(d)$, we show optical properties measured for this system, namely a transmission plot and the averaged relative change of LDOS for vector waves. We report an isotropic transmission gap, that is correlated to a depletion of DOS, which shows that $3d$ gyromorphs are promising bandgap formers.
Figure 6: Polygyromorphs. $(a)$ Structure factor featuring $3$ rings with $82$, $106$, and $134$ peaks at $KL/2\pi$ equal to $85$, $92.5$, and $100$ respectively. Corresponding $(b)$ $g(\bm{r})$, $(c)$ TE transmission, and $(d)$ Radially averaged transmission $T$ (top) and LDOS relative change (bottom).
Beyond one ring – Finally, we extend the concept of gyromorphs to disordered systems displaying several rings of peaks in $S(k)$. An example system is shown in Fig. 6, where we impose $3$ rings of mutually prime (up to parity) rotational orders, at distances with irrational ratios to ensure that the structure is not trivially compatible with a single (quasi)crystal. The resulting structure factor and pair correlation function are shown in Fig. 6$(a)$ and $(b)$, respectively. This structure is expected to display $3$ consecutive dips in transmission and DOS, which we indeed report in Fig. 6$(c)$ and $(d)$. This example highlights that gyromorphs are much more general than “dual quasicrystals” and can be used to obtain arbitrary transmission gaps in an aperiodic medium.
Conclusion – We introduced a novel class of disordered materials, termed gyromorphs, which are translationally disordered but exhibit one (or several) ring of delta peaks, resulting in quasi-long-range discrete rotational order. We showed that gyromorphs with high rotational order exhibit deeper isotropic bandgaps than systems previously considered in the literature, in both 2d and 3d, and that several such features can be imposed at once. This advance is significant both fundamentally and practically. Fundamentally, we conceived and verified the existence of a new class of structures that reconciles seemingly contradictory features and stands apart from all known material types. Practically, our numerical predictions indicate that gyromorphs outperform existing isotropic photonic bandgap materials. Gyromorphs thus show promise for applications such as freeform waveguides [1, 2] and lightweight coatings with strong backscattering, such as highly reflective “white” structures [59, 60].
Acknowledgements.
Acknowledgments – The authors would like to thank Weining Man, Paul Chaikin, Dov Levine, Rémi Carminati, Luís Froufe-Pérez, Frank Scheffold, and Werner Krauth for insightful comments on this work. A.S., M.C., and S.M. acknowledge the Simons Center for Computational Physical Chemistry for financial support. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. Parts of the results in this work make use of the colormaps in the CMasher package [61]. Correlations are computed using the rusted_core library. The FReSCo, MAGreeTe, and rusted-core libraries are publically available [62, 51, 63].